OpenStudy (anonymous):

Your piggy bank has a total of 47 coins in it; some are dimes and some are nickels. If you have a total of $3.95, how many nickels do you have? @JakeV8

OpenStudy (anonymous):

dimes are 10 cents each and nickels are 5 cents each right?

OpenStudy (anonymous):

the first two equations would be n+d=47 & .05n+.10d=3.95 but idk how to get the rest

OpenStudy (anonymous):

solve them by elimination or substitution

OpenStudy (anonymous):

substitution : n = 47 - d , then substitute this n into the other equation. elimination : .05 n + .1 d = 3.95 times 20 = n + 2d = 79.0 n + 2d = 79 n + d = 47 minus them :x

OpenStudy (anonymous):

i dnt understand...

OpenStudy (anonymous):

n + d = 47 --> n = 47 - d then put n above into 0.05n + 0.1d = 3.95

OpenStudy (anonymous):

.05(47-d)+0.1d=3.95 2.35-.05d+0.1d=3.95 2.35+.05d=3.95 -2.35 -2.35 .05d=1.6 /.05 /.05 d=32 ???? did i do it correctly? then it would be 47-32=15 n=15 ??

OpenStudy (anonymous):

@nphuongsun93

OpenStudy (anonymous):

yup that's right

OpenStudy (anonymous):

thx!